Following the Task 1 experimental procedure as per the Project 9 Guide (pages 5 – 7), 2.006 g of coconut oil is run and 0.796 g of fatty acid salts (soap) is obtained. Calculate the percent yield. Show all work. % yield= (actual yield)/(theoretical yeild )(100) Oil MM = 869.4 g/mol Sodium Fatty Acid Salt (Soap): 299.1 g/mol Potassium Fatty Acid Salt (Soap): 315.2 g/mol
Answer #1

Answer:

Given,

To determine percentage yield

By using the given molar masses,

869.4 g of olil = (299.1+315.2) g Fatty acid salt.

I.e.,

869.4 g of olil = 614.3 g Fatty acid salt.

Theoretical Yield:

If 0 869.4 g of olil = 614.3 g Fatty acid salt.

Then 2.006 g of olil = say "A" g Fatty acid salt

A x 869.4 = 2.006 x 614.3

A = 2.006 x 614.3? / 869.4.

A = 1.4174 g

Theoretical Yield of Fatty Acid = 1.4174 g.

Now,

Actual yield

Actual yield of Fatty Acid = 0.796 g

% Yield = (0.796/1.4174) x100.

% yield = 0.562*100

% Yield = 56.2%